# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ Abstract
We study the task of learning Multi-Index Models (MIMs) in the presence of label noise under the Gaussian distribution. A $K$-MIM on $\mathbb{R}^{d}$ is any function $f$ that only depends on a $K$-dimensional subspace, i.e., $f(\mathbf{x})=g(\mathbf{W}\mathbf{x})$ for a link function $g$ on $\mathbb{R}^{K}$ and a $K\times d$ matrix $\mathbf{W}$. We consider a class of well-behaved MIMs with finite ranges that satisfy certain regularity properties. Our main contribution is a general noise-tolerant learning algorithm for this class that is qualitatively optimal in the Statistical Query (SQ) model. At a high-level, our algorithm attempts to iteratively construct better approximations to the defining subspace by computing low-degree moments of our function conditional on its projection to the subspace computed thus far, and adding directions with relatively large empirical moments. For our class of well-behaved MIMs, we show that this procedure efficiently finds a subspace $V$ so that $f(\mathbf{x})$ is close to a function of the projection of $\mathbf{x}$ onto $V$, which can then be found by brute-force. Conversely, for functions for which these conditional moments do not necessarily help in finding better subspaces, we prove an SQ lower bound providing evidence that no efficient algorithm exists.
@@ 1 Introduction
Motivation: Multiclass Linear Classification.
@@
A multiclass linear classifier is a function $f:\mathbb{R}^{d}\to[K]$ of the form $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})\;,$ where $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ and $t_{i}\in\mathbb{R}$. Where ties are resolved by selecting the smallest index. Multiclass linear classification (MLC) is a textbook machine learning problem [SB14], that has also been extensively studied in the online learning literature; see, e.g., [KSST08] and references therein. In the realizable PAC setting (i.e., with clean labels), this task is efficiently solvable to 0-1 error $\epsilon$ with sample size $O(dK/\epsilon)$ via the multiclass Perceptron algorithm. On the other hand, despite extensive effort, see, e.g., [NDRT13, WLT18, LWS18, SKP+20, GWBL20], our understanding of the computational complexity of learning is poor in the noisy setting—even for random label noise.
@@
The special case of binary classification of Linear Threshold Functions (LTFs) or halfspaces—corresponding to $K=2$—is fairly well-understood. Specifically, sample and computationally efficient learning algorithms are known, even in the presence of noisy labels [ABL17, DKS18, DGT19, DKTZ20a, DKTZ20b, DKK+21b, DKK+22]. Unfortunately, these prior approaches are specific to the binary case. In particular, it is not clear whether they can be generalized to give non-trivial efficient algorithms for the noisy multiclass setting.
@@
Perhaps the simplest model of random label noise is that of Random Classification Noise (RCN) [AL88]. In the binary classification setting, the RCN model posits that the label of each example is independently flipped with probability exactly $\eta$, for some noise parameter $20<\eta<1/$ 2. A slight generalization of the noise model, that is essentially equivalent, allows for different probabilities $\eta_{+},\eta_{-}\in(0,1/2)$ of flipping the label of positive and negative examples. A classical result [Kea98] shows that, in the binary classification setting, any efficient Statistical Query (SQ) algorithm can be transformed into an efficient RCN-tolerant PAC learner (with a polynomial dependence in all relevant parameters). Unfortunately, no such result is known for the multiclass setting.
@@
The natural (and well-studied) generalization of the RCN model to the multiclass setting is the following [PRKM+17, ZLA21, VRW17]. Let $D$ be the distribution on noisy samples $(\mathbf{x},y)\in\mathbb{R}^{d}\times[K]$, where in our setting the marginal distribution $D_{\mathbf{x}}$ is the standard Gaussian, and $f:\mathbb{R}^{d}\to[K]$ be the ground truth classifier. Then there exists a row-stochastic matrix $\mathbf{H}\in\mathbb{R}_{+}^{K\times K}$ such that $\mathbf{H}_{i,j}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid f(\mathbf{x})=j]$. Moreover, it is additionally assumed that there exists $\gamma>0$ such that $\mathbf{H}_{i,i}\geq\mathbf{H}_{i,j}+\gamma$, for all $i\neq j\in[K]$ [^1] Roughly speaking, the parameter $\gamma>0$ plays the role of $1-2\eta$ in the binary case.. Under these assumptions, MLC is information-theoretically solvable using $O(dK/(\gamma\epsilon))$ samples. On the other hand, no $\mathrm{poly}(d,K)$ time learning algorithm is known. Prior algorithmic work using gradient-based methods to achieve sample and time complexity that scales polynomially with $1/\sigma$, where $\sigma$ is the minimum singular value of $\mathbf{H}$ (see [LWS18]). This quantity could be arbitrarily small or even zero, hence such approaches do not in general lead to efficient algorithms. This discussion leads to the following question:
@@
A related question is to understand the algorithmic possibilities and limitations of learning MLCs in the presence of adversarial label noise. In the adversarial label noise model, an adversary is allowed to arbitrarily corrupt an unknown $\mathrm{OPT}$ fraction of the labels, for some parameter $\mathrm{OPT}<1/2$, and the goal of the learner is to compute a hypothesis whose 0-1 error is competitive with $\mathrm{OPT}$ (see Definition 1.3). The information-theoretically optimal error guarantee in this model is $\mathrm{OPT}+\epsilon$. Unfortunately, even for $K=2$, known computational lower bounds rule out the existence of such algorithms with complexity better than $d^{\mathrm{poly}(1/\epsilon)}$ [DKPZ21, DKR23]. To achieve a fixed polynomial runtime in $d$, one needs to relax the error guarantee. In particular, for $K=2$, efficient algorithms are known with complexity $\mathrm{poly}(d/\epsilon)$ and error guarantee $O(\mathrm{OPT})+\epsilon$. A natural question is whether qualitatively similar algorithms exist for the multiclass setting. Specifically, a first goal is to establish a fixed-parameter tractability guarantee:
@@
The link function $g$ could be either a priori known or unknown to the learner. In the latter case, it is typically assumed to satisfy certain regularity conditions. Moreover, the dimension $K$ of the hidden subspace is usually significantly smaller than the ambient dimension $d$.
@@ 1.1.1 Learning Well-behaved Multi-Index Models with Noise
Our general algorithmic result applies to MIMs with finite range that satisfy appropriate regularity properties. Throughout this section, we will use $d$ for the dimension of the feature space and $K$ for the dimension of the hidden low-dimensional subspace. A standard assumption in the literature on MIMs, satisfied by a wide range of natural concept classes, is that the underlying function has bounded Gaussian surface area (GSA)—a complexity measure of the decision boundary (see Definition 3.1 for the definition in the multiclass case). We denote by $\Gamma$ the upper bound on the GSA of the underlying concept class.
The key structural property that we will assume for every MIM function $f$ in our concept class is, roughly speaking, the following: it is not possible to add noise to $f$ to conceal crucial directions so that they remain invisible to moment-based detection. Specifically, for any subspace $V$, one of the following must hold:
@@
    Otherwise, there is at least one point in $V$ such that when we condition on $\mathbf{x}^{V}$ near that point, the conditional distribution of a “noisy” version of $f$ exhibits a distinguishing moment of degree at most $m$. That is, there exists a low-degree polynomial whose conditional expectation, conditioned on $\mathbf{x}^{V}$, cannot remain hidden; and hence reveals a dependence on directions outside the subspace $V$.
@@
Fix $m\in\mathbb{Z}_{+},\zeta,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality. We say that a $K$-MIM $f:\mathbb{R}^{d}\to\mathcal{Y}$ with GSA at most $\Gamma$ is is $(m,\zeta,\tau,\Gamma)$-well-behaved if for any random variable $y(\mathbf{x})$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq y(\mathbf{x})]\leq\zeta$ and any subspace $V\subseteq\mathbb{R}^{d}$, one of the following conditions holds:
@@
    where $U=(V+W)\cap V^{\perp}$ and $W$ is the hidden subspace defining $f$.
@@
Let $\mathcal{Y}$ be finite and $\cal{C}$ a class of functions from $\mathbb{R}^{d}$ to $\mathcal{Y}$. Given $\epsilon>0$ and i.i.d. samples $(\mathbf{x},y)$ from a distribution $D$ on $\mathbb{R}^{d}\times\mathcal{Y}$ such that the marginal $D_{\mathbf{x}}$ is the standard normal and no assumptions are made on the labels, the goal is to output a hypothesis $h:\mathbb{R}^{d}\to\mathcal{Y}$ that, with high probability, has small 0-1 error $\mathrm{err}_{0-1}^{D}(h)\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$, as compared to the optimal 0-1 error of the class ${\cal C}$, i.e., $\mathrm{OPT}\coloneqq\inf_{c\in\mathcal{C}}\mathrm{err}_{0-1}^{D}(c)$.
@@
At a high-level, the main advantage of the above algorithm is that—as long as $m=O(1)$—its complexity is bounded by a constant-degree polynomial in $d$. Such a bound can be viewed as a significant qualitative improvement over $d^{K}$, which may be attainable using simpler methods. In other words, we give a fixed-parameter tractability result for the task of learning well-behaved MIMs. While the complexity guarantee we obtain is new even in the realizable setting, our algorithm is additionally robust to adversarial label noise—sometimes achieving near-optimal noise-tolerance (within constant factors). In addition to its own merits, the significance of our general algorithm is illustrated by its downstream applications (given in the following subsection).
@@
Theorem 1.5 (SQ Lower Bound for Learning Not-Well-behaved MIMs).
Suppose that a function class $\mathcal{G}$ is rotationally invariant and contains a bounded surface area MIM not contained in the function class $\mathcal{F}$ of Definition 1.2 for parameters $m,\tau$. Then, any SQ algorithm that learns $\mathcal{G}$ within error smaller than $O(\tau)$ has complexity at least $d^{\Omega(m)}$.
@@ 1.1.2 Concrete Learning Applications to Well-studied Concept Classes
Multiclass Linear Classifiers.
@@
Theorem 1.6 is obtained from Theorem 1.4 via a new structural result that we establish, roughly showing that the class $\mathcal{L}_{d,K}$ is $(1,\zeta,O(\zeta),O(K))$-well behaved for any $\zeta\in(0,1)$ (see Proposition 2.4).
Theorem 1.6 gives the first constant-factor approximate learner for multiclass linear classification with adversarially corrupted labels, whose complexity is a fixed-degree polynomial in $d$. We provide some remarks to interpret the statement of Theorem 1.6. First, even for $K=2$, known hardness results imply that achieving error $\mathrm{OPT}+\epsilon$ requires $d^{\mathrm{poly}(1/\epsilon)}$ time [DKPZ21, DKR23]. Hence, to achieve a fixed-degree $\mathrm{poly}(d)$ complexity guarantee, it is necessary to relax the final accuracy. The best thing to hope for here is a constant-factor approximation, i.e., a final error of $O(\mathrm{OPT})+\epsilon$ (like the one we achieve). For the $K=2$ case, prior work achieves such an error bound [DKS18] [^3] Earlier work [ABL17] achieves such a constant-factor approximation for the special case of homogeneous halfspaces.; and in fact with complexity $\mathrm{poly}(d/\epsilon)$. On the other hand, no non-trivial approximation algorithm was previously known for the multiclass case. Importantly, no polynomial-time learner with error $o(d)\,\mathrm{OPT}+\epsilon$ is known even for $K=3.$ The crucial difference is that the approximation ratio of our algorithm is an absolute constant, independent of the dimension and the number of classes.
@@
Interestingly, using a variant of our general MIM algorithm, we obtain an RCN-tolerant learner with $O(\mathrm{OPT})+\epsilon$ error, whose complexity is qualitatively better than our agnostic learner. Specifically, we show the following (see Theorem 4.9 for the detailed statement):
@@
Intersections of Halfspaces.
@@
Prior work [KKMS05, KOS08] demonstrated that the class $\mathcal{F}_{d,K}$ is agnostically learnable under the Gaussian distribution, within optimal error of $\mathrm{OPT}+\epsilon$, using $d^{O({\log K}/{\epsilon^{4}})}$ samples and runtime. This complexity bound was recently shown to be near-optimal in the SQ model [DKPZ21, HSSV22]. A line of subsequent work has focused on understanding the tradeoff between accuracy and computational complexity, and in particular on designing faster algorithms with relaxed error guarantees. Specifically, [DKS18] developed a $\mathrm{poly}(d)(1/\epsilon)^{\mathrm{poly}(K)}$ time algorithm with error of $\mathrm{poly}(K)\tilde{O}(\mathrm{OPT}^{1/11})$. Our algorithm achieves a different tradeoff: its error guarantee is significantly better compared to [DKS18]—namely nearly-optimal as a function of $\mathrm{OPT}$—while its computational complexity incurs a $2^{\mathrm{poly}(K/\epsilon)}$ dependence (as opposed to $(1/\epsilon)^{\mathrm{poly}(K)})$).

### v2

@@ Abstract
We study the task of learning Multi-Index Models (MIMs) in the presence of label noise under the Gaussian distribution. A $K$-MIM on $\mathbb{R}^{d}$ is any function $f$ that only depends on a $K$-dimensional subspace, i.e., $f(\mathbf{x})=g(\mathbf{W}\mathbf{x})$ for a link function $g$ on $\mathbb{R}^{K}$ and a $K\times d$ matrix $\mathbf{W}$. We consider a class of well-behaved MIMs with finite ranges that satisfy certain regularity properties. Our main contribution is a general noise-tolerant learning algorithm for this class whose complexity is qualitatively optimal in the Statistical Query (SQ) model. At a high-level, our algorithm attempts to iteratively construct better approximations to the defining subspace by computing low-degree moments of our function conditional on its projection to the subspace computed thus far, and adding directions with relatively large empirical moments. For well-behaved MIMs, we show that this procedure efficiently finds a subspace $V$ so that $f(\mathbf{x})$ is close to a function of the projection of $\mathbf{x}$ onto $V$, which can then be found by brute-force. Conversely, for functions for which these conditional moments do not necessarily help in finding better subspaces, we prove an SQ lower bound providing evidence that no efficient algorithm exists.
@@ 1 Introduction
Motivation: Multiclass Linear Classification
@@
A multiclass linear classifier is a function $f:\mathbb{R}^{d}\to[K]$ of the form $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})\;,$ where $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ and $t_{i}\in\mathbb{R}$ (ties are resolved by selecting the smallest index). Multiclass linear classification (MLC) is a textbook machine learning problem [SB14], that has also been extensively studied in the online learning literature; see, e.g., [KSST08] and references therein. In the realizable PAC setting (i.e., with clean labels), this task is efficiently solvable to any desired error $\epsilon$ with sample size $O(dK/\epsilon)$ via linear programming (or via the multiclass Perceptron algorithm for well-behaved marginal distributions, including the Gaussian). On the other hand, despite extensive effort, see, e.g., [NDRT13, WLT18, LWS18, SKP+20, GWBL20], our understanding of the computational complexity of learning is poor in the noisy setting—even for the simplest case of random label noise.
@@
The special case of binary classification of Linear Threshold Functions (LTFs) or halfspaces—corresponding to $K=2$—is fairly well-understood. Specifically, sample and computationally efficient learning algorithms are known, even in the presence of partially corrupted labels [ABL17, DKS18, DGT19, DKTZ20a, DKTZ20b, DKK+21b, DKK+22]. Unfortunately, these prior approaches are specific to the binary case. In particular, it is unclear whether they can be generalized to give any non-trivial guarantees for the noisy multiclass setting.
@@
Perhaps the simplest model of random label noise is that of Random Classification Noise (RCN) [AL88]. In the binary classification setting, the RCN model posits that the label of each example is independently flipped with probability exactly $\eta$, for some noise parameter $20<\eta<1/$ 2. A slight generalization of the noise model, that is essentially equivalent, allows for different probabilities $\eta_{+},\eta_{-}\in(0,1/2)$ of flipping the label of positive and negative examples. A classical result [Kea98] shows that, in the binary classification setting, any efficient Statistical Query (SQ) algorithm can be transformed into an efficient RCN-tolerant PAC learner (with a polynomial dependence in all relevant parameters). Unfortunately, no such result is known for the multiclass setting [^1] In fact, Theorem 1.9 implies that such a generalization is impossible without strong assumptions on the noise matrix..
@@
The standard generalization of the RCN model to the multiclass setting is the following [PRKM+17, ZLA21, VRW17]. Let $D$ be the distribution on noisy samples $(\mathbf{x},y)\in\mathbb{R}^{d}\times[K]$, where in our setting the marginal distribution $D_{\mathbf{x}}$ is the standard Gaussian, and $f:\mathbb{R}^{d}\to[K]$ be the ground truth classifier. Then there exists a row-stochastic matrix $\mathbf{H}\in\mathbb{R}_{+}^{K\times K}$ such that $\mathbf{H}_{i,j}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid f(\mathbf{x})=j]$. Moreover, it is additionally assumed that there exists $\gamma>0$ such that $\mathbf{H}_{i,i}\geq\mathbf{H}_{i,j}+\gamma$, for all $i\neq j\in[K]$ [^2] Roughly speaking, the parameter $\gamma>0$ plays the role of $1-2\eta$ in the binary case.. Under these assumptions, MLC is information-theoretically solvable using $O(dK/(\gamma\epsilon))$ samples. On the other hand, no $\mathrm{poly}(d,K)$ time learning algorithm is known. Prior algorithmic work using gradient-based methods to achieve sample and time complexity that scales polynomially with $1/\sigma$, where $\sigma$ is the minimum singular value of $\mathbf{H}$ (see [LWS18]). This quantity could be arbitrarily small or even zero, hence such approaches do not in general lead to a computationally efficient algorithm. This discussion leads to the following question:
@@
An even more challenging goal is to understand the algorithmic possibilities and limitations of learning MLCs in the presence of adversarial label noise. In the adversarial label noise model, an adversary is allowed to arbitrarily corrupt an unknown $\mathrm{OPT}$ fraction of the labels, for some parameter $\mathrm{OPT}<1/2$, and the goal of the learner is to compute a hypothesis whose 0-1 error is competitive with $\mathrm{OPT}$ (see Definition 1.4). The information-theoretically optimal error guarantee in this model is $\mathrm{OPT}+\epsilon$. Unfortunately, even for $K=2$, known computational lower bounds rule out the existence of such algorithms with complexity better than $d^{\mathrm{poly}(1/\epsilon)}$ [DKPZ21, DKR23].
Consequently, to achieve a fixed polynomial runtime in $d$, one needs to relax the error guarantee. For the special case $K=2$, such algorithms are known. In particular, [DKS18] gave a $\mathrm{poly}(d/\epsilon)$ time algorithm with error guarantee $C\cdot\mathrm{OPT}+\epsilon$, where $C$ is a universal constant.
A natural question is whether qualitatively similar algorithms exist for the multiclass setting. Specifically, a first goal is to establish a fixed-parameter tractability guarantee:
@@
In typical applications, the dimension $K$ of the hidden subspace is usually significantly smaller than the ambient dimension $d$.
In regards to the problem of learning MIMs from labeled examples, there two qualitatively different versions: the link function $g$ in Definition 1.1 could be a priori known or unknown to the learner. In the case of unknown $g$, it is typically assumed that it satisfies certain regularity conditions; as without any assumptions on $g$, learning is information-theoretically impossible. In this work, we focus on the setting where $g$ is unknown.
@@ 1.1 Our Results
In this section, we present our main results. For a concise overview of our algorithmic contributions, see Table 1.
Function Class Prior Work Our Work
Runtime Error Runtime Error
[…]
$K$-MLC with RCN $d^{\mathrm{poly}(K/\epsilon)}$ $\mathrm{OPT}+\epsilon$ (Theorem 1.11) $\mathrm{poly}(d)(1/\epsilon)^{\mathrm{poly}(K)}$ $O(\mathrm{OPT})+\epsilon$
[…]
Table 1: Comparison of runtime and error guarantees between state-of-the-art efficient algorithms (see [KOS08, DKS18] for intersections of $K$ halfspaces) and our subspace approximation procedure.
@@ 1.1.1 Learning Well-behaved Multi-Index Models with Noise
Our general algorithmic result (Theorem 1.5) applies to MIMs with finite range that satisfy a collection of regularity properties (Definition 1.2).
Throughout this section, we will use $d$ for the dimension of the feature space and $K$ for the dimension of the hidden low-dimensional subspace. A standard assumption in the literature on MIMs (see, e.g., [KOS08, DKK+23]), satisfied by a wide range of natural concept classes, is that the underlying function has bounded Gaussian surface area (GSA)—a complexity measure of the decision boundary (see Definition 3.1 for the definition in the multiclass case). We denote by $\Gamma$ the upper bound on the GSA of the underlying concept class.
The structural property that we will assume for every MIM function $f$ in our class is, roughly speaking, the following: it is not possible to add noise to $f$ to conceal crucial directions so that they remain invisible to moment-based detection. Specifically, for any subspace $V$, one of the following must hold:
@@
    Otherwise, there is at least one point in $V$ such that when we condition on $\mathbf{x}^{V}$ near that point, the conditional distribution of a “noisy” version of $f$ exhibits a distinguishing moment of degree at most $m$. That is, there exists a degree-$m$ polynomial whose conditional expectation, conditioned on $\mathbf{x}^{V}$ (the projection of $\mathbf{x}$ on $V$), cannot remain hidden; and hence reveals a dependence on directions outside the subspace $V$.
The parameter $m$ above turns out to essentially characterize the complexity of the learning problem. We show, qualitatively speaking, that there exists an algorithm with complexity $d^{m}$; and that this bound is essentially best possible in the SQ model.
@@
Fix $m\in\mathbb{Z}_{+},\zeta,\tau\in(0,1)$, $\Gamma>0$, and $\mathcal{Y}\subseteq\mathbb{Z}$ a set of finite cardinality. We say that a $K$-MIM $f:\mathbb{R}^{d}\to\mathcal{Y}$ is $(m,\zeta,\tau,\Gamma)$-well-behaved if $f$ has Gaussian surface area at most $\Gamma$, and for any random variable $y(\mathbf{x})$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq y(\mathbf{x})]\leq\zeta$ and any subspace $V\subseteq\mathbb{R}^{d}$, one of the following conditions holds:
@@
    where $U=(V+W)\cap V^{\perp}$ is the projection of $W$ onto $V^{\perp}$ and $W$ is the subspace defining $f$.
@@
Remark 1.3.
Definition 1.2 implicitly introduces a complexity measure for learning a family of discrete MIMs. Let $\mathcal{C}$ be a family of MIMs with Gaussian Surface Area bounded by $\Gamma$, and let $\zeta,\tau>0$ be parameters. Define $m^{\ast}$ to be the smallest integer for which every $f\in\mathcal{C}$ is $(m^{\ast},\zeta,\tau,\Gamma)$-well-behaved. It turns out that the quantity $m^{\ast}$ is (essentially) a generalization of the generative exponent [DPVLB24], a measure characterizing the SQ complexity of learning Single-Index Models (SIMs). For further details, see Appendix D.
@@
Let $\mathcal{Y}$ be finite and $\cal{C}$ a class of functions from $\mathbb{R}^{d}$ to $\mathcal{Y}$. Given i.i.d. samples $(\mathbf{x},y)$ from a distribution $D$ on $\mathbb{R}^{d}\times\mathcal{Y}$ such that the marginal $D_{\mathbf{x}}$ is the standard normal and no assumptions are made on the labels, the goal is to output a hypothesis $h:\mathbb{R}^{d}\to\mathcal{Y}$ that, with high probability, has small 0-1 error $\mathrm{err}_{0-1}^{D}(h)\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$, as compared to the optimal 0-1 error of the class ${\cal C}$, i.e., $\mathrm{OPT}\coloneqq\inf_{c\in\mathcal{C}}\mathrm{err}_{0-1}^{D}(c)$.
@@
At a high-level, the main advantage of the above algorithm is that—as long as $m=O(1)$—its complexity is bounded by a constant-degree polynomial in $d$. In other words, we give a fixed-parameter tractability result for the task of learning well-behaved MIMs. While the complexity guarantee we obtain is new even in the realizable setting, our algorithm is additionally robust to adversarial label noise—sometimes achieving near-optimal noise-tolerance (within constant factors). In addition to its own merits, the significance of our general algorithm is illustrated by its downstream applications (given in the following subsection).
@@
Theorem 1.6 (SQ Lower Bound for Learning Not-Well-behaved MIMs, Informal).
Suppose that a function class $\mathcal{G}$ is rotationally invariant and contains a bounded surface area MIM not contained in the function class $\mathcal{F}$ of Definition 1.2 for parameters $m,\tau,\zeta$. Then, any SQ algorithm that learns $\mathcal{G}$ in the presence of $\zeta$ label noise to error smaller than $\tau-O(\zeta)$ has complexity at least $d^{\Omega(m)}$.
Remark 1.7 (Characterization).
Theorems 1.5 and 1.6 together essentially characterize the complexity of learning a MIM family. In particular, let $\mathcal{C}$ be a rotationally invariant family of MIMs with Gaussian Surface Area bounded by $\Gamma$, and let $\zeta,\tau>0$ be parameters. Define $m^{\ast}$ to be the smallest integer for which every $f\in\mathcal{C}$ is $(m^{\ast},\zeta,\tau,\Gamma)$-well-behaved.
Then there is an SQ algorithm that learns a (slightly less than) $\zeta$-noisy function from $\mathcal{C}$ to error $\tau+O(\zeta)$ with resources $d^{m^{\ast}}$ (times some function of the other parameters), but not an SQ algorithm that learns a $\zeta$-noisy function from $\mathcal{C}$ to error $\tau-O(\zeta)$ with resources $d^{o(m^{\ast})}$.
@@ 1.1.2 Concrete Learning Applications to Well-studied Concept Classes
Multiclass Linear Classifiers
@@
Theorem 1.8 can be derived from Theorem 1.5 via a new structural result that we establish, roughly showing that the class $\mathcal{L}_{d,K}$ is $(1,\zeta,O(\zeta),O(K))$-well behaved for any $\zeta\in(0,1)$ (see Proposition 2.4). For the sake of exposition, in Section 2 we give a direct algorithm establishing Theorem 1.8.
Theorem 1.8 gives the first constant-factor approximate learner for multiclass linear classification with adversarially corrupted labels, whose complexity is a fixed-degree polynomial in $d$. We provide some remarks to interpret this statement. First, even for $K=2$, known hardness results imply that achieving error $\mathrm{OPT}+\epsilon$ requires $d^{\mathrm{poly}(1/\epsilon)}$ time [DKPZ21, DKR23]. Hence, to achieve a fixed-degree $\mathrm{poly}(d)$ complexity guarantee, it is necessary to relax the final accuracy. The best thing to hope for here is a constant-factor approximation, i.e., a final error of $O(\mathrm{OPT})+\epsilon$ (like the one we achieve). For the $K=2$ case, prior work achieves such an error bound [DKS18] [^4] Earlier work [ABL17] achieves such a constant-factor approximation for the special case of homogeneous halfspaces.; and in fact with complexity $\mathrm{poly}(d/\epsilon)$. On the other hand, no non-trivial approximation algorithm was previously known for the multiclass case. Importantly, no polynomial-time learner with error $o(d)\,\mathrm{OPT}+\epsilon$ is known even for $K=3.$ The crucial difference is that the approximation ratio of our algorithm is an absolute constant, independent of the dimension and the number of classes.
@@
Digression: SQ Lower Bound for MLC with Partial Labels
As an additional implication of our SQ lower bound technique establishing Theorem 1.9, we establish the first super-polynomial SQ lower bound for the well-studied task of learning multiclass linear classifiers with partial (or coarse) labels. This is a classical and widely studied model of misspecification in multi-class classification [CST11, LD14, VRW17, FKKT21]. In this model, instead of observing instance-label pairs $(\mathbf{x}^{(i)},f(\mathbf{x}^{(i)}))_{i=1}^{m}$, we observe instance-label subsets $(\mathbf{x}^{(i)},S_{i})_{i=1}^{m}$, where each $S_{i}\subseteq\mathcal{Y}$ and the true label always belongs to the observed subset (i.e., $f(\mathbf{x}^{(i)})\in S_{i}$). The goal is to find a hypothesis with error $\epsilon$ with respect to the original labels. A necessary assumption for identifiability in this model is that any incorrect label does not appear too frequently in the observed subset (as quantified by the parameter $\gamma$ in the theorem statement below); see Definition E.1 for a formal definition. Under this assumption, the sample complexity of ERM for this task is $O(dK/(\gamma\epsilon))$ [LD14]. On the computational front, we establish the following SQ lower bound (see Theorem E.2):
Theorem 1.10 (SQ Lower Bound for MLC with Partial Labels).
Any SQ algorithm that learns $\mathcal{L}_{d,K}$ given partial labels, with $\gamma=\mathrm{poly}(1/K)$, under the Gaussian distribution to error $O(1/K)$, has complexity $d^{\Omega(K)}$.
The details are deferred to Appendix E.
We now return to our algorithmic contributions. Interestingly, using a variant of our general MIM algorithm, we obtain an RCN-tolerant learner with $O(\mathrm{OPT})+\epsilon$ error, whose complexity is qualitatively better than our agnostic learner. Specifically, we show the following (see Theorem 4.9):
@@
Remark 1.12.
Recent work [DMRT25] has shown that the above error guarantee is not possible for efficient SQ algorithms in the distribution-free setting. Specifically, any SQ algorithm for MLC with RCN that achieves any constant factor approximation requires super-polynomial in $d$ time. This hardness result further motivates the study of distribution-specific learning in the multiclass setting, even with random noise only.
Intersections of Halfspaces
@@
To highlight the significance of Theorem 1.13, some remarks are in order. Note that the information-theoretically optimal error in our setting is $\mathrm{OPT}+\epsilon$. The only known algorithm achieving this guarantee is the $L_{1}$ polynomial regression [KKMS05, KOS08], which has complexity $d^{O({\log K}/{\epsilon^{4}})}$ for the class $\mathcal{F}_{d,K}$. (This complexity bound was recently shown to be near-optimal in the SQ model [DKPZ21, HSSV22].)
A fundamental open question in the literature has been whether one can achieve near-optimal error guarantees with qualitatively improved sample and computational complexity. Theorem 1.13 answers this question in the affirmative. Specifically, for $K=O(1)$, it achieves near-optimal error guarantee (up to a $O(\log(1/\mathrm{OPT}))$ factor), and its complexity has fixed-polynomial dependence in the dimension (namely, $d^{2}$ as opposed to $d^{f(k,1/\epsilon)}$).
In comparison, [DKS18] developed a $\mathrm{poly}(d)(1/\epsilon)^{\mathrm{poly}(K)}$ time algorithm, achieving an error guarantee of $\mathrm{poly}(K)\tilde{O}(\mathrm{OPT}^{1/11})$. The key distinction is that the error guarantee achieved by the [DKS18] algorithm is highly suboptimal as a function of $\mathrm{OPT}$ (by a large polynomial factor). Additionally, the dependence on $K$ in our Theorem 1.13 is significantly better by a polynomial factor.